OpenStudy (anonymous):

find the x and y intercept of f(x)=3x-6 and how would i graph the answer

Directrix (directrix):

To get the x-intercept, let y = 0. y= 3x - 6 0 = 3x - 6 -3x = -6 x = 2 or (2,0) y-intercept Let x = 0. y= 3x - 6 y = 3(0) - 6 y = 6 or (0,6) Plot the two points and draw the line joining them.

Directrix (directrix):